A runner completes the first 10 laps of a 15 - lap race at an average speed of 1 minute and 30 seconds per lap. How fast will the runner’s average lap time need to be for the last five laps if they want to have average per-lap speed of 1 minute and 25 seconds for the entire race?
Answer #1

here average time for first 10 laps =1.5 minutes =90 seconds

let average time for last 5 laps = a seconds/lap

therefore weighted average lap time =(10*90+a*5)/15 =85 seconds

900+5a =1275

a =(1275-900)/5= 75 seconds or 1 minute and 15 seconds